OpenStudy (hawk):

Ok let me make a easy question FIND d^2y/dx^2 if y=log(x^2/e^x)

OpenStudy (agent0smith):

First time you've used parentheses! Except we don't know what base the log is.

OpenStudy (hawk):

\[\log x^{2}\div e^{x}\]

OpenStudy (agent0smith):

Now it's even less clear than it was.

OpenStudy (hawk):

\[\frac{ x^{2} }{ e^{x} }\]

OpenStudy (agent0smith):

Is it log base 10 I guess??

OpenStudy (hawk):

\[\log \frac{ x ^{2} }{ e ^{x} }\]

OpenStudy (hawk):

Thats super clear

OpenStudy (agent0smith):

So it's log base 10, correct? Then you may want to use the change of base formula first, get it to base e.

OpenStudy (hawk):

IDK

OpenStudy (hawk):

thats why i am asking

OpenStudy (agent0smith):

Uhhh if you don't know if it's log base 10, how am i supposed to?

OpenStudy (hawk):

It is not log base 10 lol use log formula

OpenStudy (agent0smith):

Then what base is it?!

OpenStudy (hawk):

log/log

OpenStudy (hawk):

Lol you can that eqn like this too

OpenStudy (hawk):

\[\log x^{2}-e^{x}\]

OpenStudy (hawk):

standard log formulas

OpenStudy (agent0smith):

Yeah, so i guess we can assume it's log base 10... so change the base to e, you should know how to do that

OpenStudy (agent0smith):

\[\large y = \frac{ \ln x^2 }{ \ln 10 }-\frac{ \ln e^x }{ \ln 10 }\] \[\large y = \frac{ 2\ln x }{ \ln 10 }-\frac{ x }{ \ln 10 }\]

OpenStudy (mww):

Assuming natural logarithm (most common anyway) \[y = \ln (\frac{ x^2 }{ e^x }) = \ln (x^2) - \ln(e^x) = \ln(x^2) - x\ = 2 \ln(x) - x\] \[\frac{ dy }{ dx } = \frac{ 2 }{ x } - 1\] \[\frac{ d^2y }{ dx^2 } = -\frac{ 2 }{ x^2 }\]